# Retrieval-Augmented Mixture of LoRA Experts for Uploadable Machine Learning

### v1

@@ Abstract.
Low-Rank Adaptation (LoRA) provides an effective yet efficient solution for fine-tuning large language models (LLMs). The modular and plug-and-play nature of LoRA enables the integration of diverse domain-specific LoRAs to enhance the capabilities of LLMs. The emergence of open-source platforms like Huggingface and Modelscope has given rise to a new computational paradigm termed Uploadable Machine Learning (UML). In this framework, contributors on the edge-side use decentralized data from various domains to train specialized adapters. These adapters are then uploaded to a central platform to improve the capabilities of LLMs. The platform leverages the plug-and-play nature of various domain-specific adapters to address downstream heterogeneous requests. These requests are often in a mixed-task format, necessitating personalized service. Previous research on composing multiple LoRAs either focuses on specific isolated downstream tasks or fixes the selection of LoRAs during training. However, in the context of UML, the pool of candidate LoRAs is not only dynamically updated but also continuously expanded with new LoRAs being uploaded to expand the capabilities of LLMs in specific domains. This requires the platform’s LoRA selection/routing mechanism to be generalizable to unseen LoRAs. Furthermore, the downstream tasks are of a mixed-task nature, requiring the provision of personalized services to accommodate the heterogeneous requests. To bridge this gap, we propose Retrieval-Augmented Mixture of LoRA Experts (RAMoLE), a retrieve-then-compose framework that adaptively retrieves and composes multiple LoRAs according to the input prompts. RAMoLE contains three main components: firstly, identifying and retrieving LoRAs relevant to the given input through LoraRetriever; secondly, coordinating various dynamically loaded LoRA modules from the retriever through a carefully designed on-the-fly MoLE mechanism; and thirdly, developing efficient batch inference to accommodate heterogeneous requests. The experimental results indicate that RAMoLE consistently outperforms the baselines, emphasizing its effectiveness and flexible scalability.
LoRA Composition, Retrieval Augmented Large Language Model, Mixture of Experts, Uploadable Machine Learning
copyright: acmlicensedjournalyear: 2024ccs: Computing methodologies Natural language generationccs: Computing methodologies Learning paradigmsccs: Computing methodologies Cooperation and coordination
1. Introduction
@@
Figure 1. Illustration of UML. a) LoRAs from various domains and tasks aimed at enhancing specific capabilities of the LLM can be uploaded to or updated to the LoRA pool. b) The multi-LoRA serving framework aims to leverage the plug-and-play nature of LoRAs to offer comprehensive services. c) The downstream tasks, presented in a mixed-task form, require personalized expert routing.
@@
Recently, leveraging the principles of scaling laws (Kaplan et al., 2020), Large Language Models (LLMs) like openAI’s GPT-3.5 (Ouyang et al., 2022), GPT-4 (Achiam et al., 2023), Meta’s LLama (Touvron et al., 2023) and many other Foundation Models (Wei et al., 2021; Chowdhery et al., 2023; Bai et al., 2023; Young et al., 2024) have achieved notable success across various natural language processing (NLP) tasks (Hadi et al., 2023; Wang et al., 2023). However, these models, primarily trained on diverse internet datasets, sometimes struggle in specialized areas. To address this, numerous studies (Liu et al., 2023; Yang et al., 2023; Hadi et al., 2023) have explored fine-tuning LLMs using domain-specific data, thereby adapting them to fulfill specific needs. Due to the prohibitively high computation costs for fine-tuning LLMs on specific domains, there is a growing shift towards Parameter-Efficient Fine-Tuning (PEFT) (Liu et al., 2022; Hu et al., 2023, 2021), which only updates a small fraction of the model’s parameters or integrates new trainable parameters that augment the model’s capabilities.
@@
As interest grows in open-source communities like Hugging Face [^1] https://huggingface.co/, ModelScope [^2] https://modelscope.cn/home, and Civit AI [^3] https://civitai.com/, there is an increasing trend toward sharing numerous PEFT modules trained on a broad spectrum of domains and tasks across these platforms. Within these PEFT methods, LoRA has become the dominant fine-tuning method (Luo et al., 2024). LoRA achieves efficiency by decomposing the weight matrices of a pre-trained model into low-rank factors, allowing it to adapt the model with minimal parameter changes. This approach not only significantly reduces computational load but also maintains or improves model performance through targeted updates. Its modularity and plug-and-play capabilities make it highly adaptable for various tasks and applications. This trend is giving rise to a new decentralized computing paradigm, which we refer to as Uploadable Machine Learning (UML) (Zhao et al., 2024; Luo et al., 2024; Liu et al., 2024). As shown in Fig.1, this new computing paradigm involves three components: a) The edge-side domain-specific LoRA contributors. These contributors train plug-and-play, domain-specific LoRA modules in a decentralized manner using local, domain-specific data. The trained LoRA parameters are then uploaded to a centralized service platform. b) Centralized Multi-LoRA Serving platform. The server side operates an LLM equipped with a dynamic pool of LoRAs. The purpose of the LoRA pool is to enhance the LLM’s capabilities across various domains, allowing it to deliver strong performance in a wider range of fields. The LLM leverages the plug-and-play capability of LoRA to dynamically load various LoRAs, allowing it to deliver personalized services tailored to diverse downstream requests. c) Mixed-task Downstream Heterougenous Requests. The downstream requests cover a diverse range of tasks and are in mixed-task form. The heterogeneous nature of these requests requires that the server side utilize the appropriate LoRA for different prompts to deliver personalized services.
@@
Recently, some work has focused on how to compose different LoRAs to enhance the model’s capabilities. Several research has explored the integration of the mixture of expert (MoE; Jacobs et al. (1991); Jordan and Jacobs (1994)) with LoRAs (Wang et al., 2022; Liu et al., 2023; Zadouri et al., 2023; Muqeeth et al., 2023; Wu et al., 2023a). However, these methods fix the selection of LoRAs during training, lacking the flexibility to dynamically update and scale in scenarios where the LoRA pool may consistently expand in the UML scenario. This rigidity violates the inherently dynamic and uploadable nature of decentralized training involving domain-specific LoRA experts. LoRAHub (Huang et al., 2023) and AdapterSoup (Chronopoulou et al., 2023) explore composing LoRAs for specific downstream tasks. However, these two methods provide a one-size-fits-all approach to downstream tasks, failing to accommodate the heterogeneous nature of diverse requests in a UML setting.
In this paper, we introduce a Retrieval-Augmented Mixture of Experts (RAMoLE), a retrieve-then-compose framework designed to exploit the plug-and-play nature of LoRA for UML. Our framework consists of three key components: (1) Input-aware LoRA Retrieval: The first step of our framework is aligning user inputs with the corresponding LoRAs through sentence embeddings and is further refined by an instruction fine-tuning  (Su et al., 2022; Asai et al., 2022) for effective LoRA retrieval. Through the retriever, we achieve a more flexible LoRA routing mechanism, whose training stage is disentangled from the training and inference of the LLM. (2) On-the-fly Mixture of LoRA Experts: After retrieving the corresponding LoRAs for each sample, we need to finely differentiate and compose these LoRAs for targeted generation. To this end, we propose a novel MoLE mechanism that involves training an additional RouterLoRA, which can be seamlessly integrated into the LLM in an off-the-shelf manner. This mechanism uses an attention mechanism to determine the weights corresponding to the retrieved LoRAs, thereby achieving a more flexible and dynamic MoLE architecture. It is worth noting that vanilla LoRA composition methods, which involve either averaging parameters or outputs, assign uniform weight to all loaded LoRAs. Such approaches may not allocate the most effective weights to the most suitable LoRA, potentially leading to suboptimal outcomes. In contrast, our proposed MoLE mechanism, utilizing a RouterLoRA, finely tunes the allocation of weights to each LoRA, dynamically differentiating each one to enhance overall performance (3) Batch Inference of Multiple LoRAs: Most previous work on the input-adaptive inference of LLMs does not support batch inference (Zhou et al., 2020; Chronopoulou et al., 2023). To tackle the challenge of heterogeneous batched requests, we construct a unique LoRA mapping matrix for batch samples. This allows for tailored inferences through efficient matrix multiplication, ensuring each request activates its corresponding LoRAs while maintaining batch processing efficiency.
@@
We note that a shorter conference version of this paper appeared in (Zhao et al., 2024). Our conference version of the paper did not explicitly define the entire scenario. This manuscript more thoroughly and comprehensively defines the setting as Uploadable Machine Learning. Additionally, the conference version simply applied the same weights for composing the retrieved LoRAs, an approach that may not allocate the most effective weights to the most appropriate LoRA, potentially resulting in suboptimal outcomes. In this manuscript, we introduce a novel on-the-fly MoLE method that employs a RouterLoRA to distinguish between each retrieved LoRA, enhancing performance in the UML scenario. Additionally, to accommodate the new MoLE mechanism, the batch inference mechanism has been adapted to facilitate a more efficient inference process.
Our contribution can be summarized as follows:
@@
2. Related Work
2.1. Mixture of Experts
@@
The Mixture of Experts (MoE) method combines various specialized sub-modules, guided by a gating network to tailor responses to different input types (Jacobs et al., 1991; Jordan and Jacobs, 1994; Shen et al., 2023; Riquelme et al., 2021; Dou et al., 2023). Some work (Wang et al., 2022; Zadouri et al., 2023; Zhu et al., 2023; Liu et al., 2023; Dou et al., 2023; Ostapenko et al., 2023; Feng et al., 2024) focuses on using the MoE method for PEFT to achieve more effective and efficient model fine-tuning. Other work (Wu et al., 2023a; Muqeeth et al., 2023) focuses on using MoE to post-hoc coordinate existing LoRA experts without specifically training the experts’ parameters. Specifically, AdaMix (Wang et al., 2022) introduces a mixture-of-adaptations approach for fine-tuning models efficiently by combining multiple adaptation methods to reduce parameter overhead. Ostapenko et al. (2023) demonstrates the efficacy of combining MoE with lightweight parameter-efficient tuning methods to enhance instruction tuning for large language models, achieving performance comparable to full fine-tuning while significantly reducing computational resource requirements. Zadouri et al. (2023) enhances MoE by developing a highly parameter-efficient model that achieves performance on par with full fine-tuning while significantly reducing parameter usage. SiRA (Zhu et al., 2023) introduces a sparse mixture approach to low-rank adaptation, improving the efficiency and performance of large language models by optimizing sparse computation and reducing overfitting. MoELoRA (Liu et al., 2023) leverages MoE and LoRA to efficiently fine-tune large language models for diverse medical tasks, addressing the issues of task variety and high computational cost. LoRAMoE (Dou et al., 2023) integrates MoE and LoRA to enhance large language models’ capability in maintaining and applying world knowledge efficiently, significantly improving model performance while reducing computational requirements. Mixture-of-LoRAs (Feng et al., 2024) presents a method that leverages a mixture of LoRA modules to enhance the multitask performance of large language models, achieving significant efficiency in fine-tuning across various tasks. These works focus on integrating MoE and LoRA for more efficient fine-tuning of LLMs and reducing computational resources during inference. Our work is orthogonal to these approaches, as we investigate post-hoc LoRA combination, aligning with recent trends in adapting a large LoRA pool for various downstream tasks.
@@
MoLE (Wu et al., 2023a) proposes an inference-time MoE method to enhance the parameter efficiency and flexibility of large language models, allowing for improved fine-tuning and adaptation across diverse tasks without significant computational overhead. SMEAR (Muqeeth et al., 2023) improves the performance of mixture-of-experts (MoE) models by using a soft merging approach for expert parameters combined with adaptive routing, avoiding the inefficiencies of discrete routing decisions and resulting in better expert specialization and overall model efficiency. PATGOOSE (Muqeeth et al., 2024) explores a framework for zero-shot learning where a model learns to dynamically route inputs among specialized experts, thereby achieving generalization to unseen tasks by leveraging the expertise of various pre-trained modules, significantly enhancing the model’s adaptability and performance in zero-shot scenarios. Ostapenko et al. (2024) presents a modular approach for LLMs by constructing and reusing a library of LoRA modules, enabling efficient fine-tuning and adaptation for multiple tasks, thereby facilitating the creation of versatile and resource-efficient LLMs that can be easily customized for different applications These methods necessitate that the router has encountered all LoRAs during training and requires the selection of LoRAs to be fixed at that time. This limitation prevents these methods from managing the continuously expanding LoRA pool in UML and from routing for unseen LoRAs.
2.2. Adapter Merging
@@
In addition to model ensembling through the MoE, there is an increasing focus on aggregating adapters from different domains through the method of Adapter Merging. Zhang et al. (2023) proposes to compose these parameter-efficient modules through linear arithmetic operations in the weight space to integrate different module capabilities. AdapterSoup (Chronopoulou et al., 2023) aggregates different adapters in the parameter space, allowing large language models to adapt to new domains without additional training. LoRAHub (Huang et al., 2023) employs random sampling of LoRA parameters from various domains and tasks, followed by black-box optimization to learn the weights of different LoRA parameters without involving model gradient backpropagation. LoRA-Flow (Wang et al., 2024) dynamically fuses multiple Low-Rank Adaptation (LoRA) modules to enhance the generative capabilities of LLMs, enabling more efficient and effective task-specific fine-tuning without substantial computational overhead. $\pi$-Tuning  (Wu et al., 2023b) considers not only language model experts but also visual model experts. It obtains weights for different LoRAs through task embeddings for downstream tasks, similarly utilizing weighted averaging in the parameter space for targeted model parameter adaptation. LoRA-Composer (Yang et al., 2024) proposes a framework that utilizes Low-Rank Adaptation to enable multi-concept customization in training-free diffusion models, addressing the challenge of integrating multiple concept LoRAs for complex and varied image generation tasks. Zhong et al. (2024) explores Multi-LoRA Composition for combining multiple LoRA modules to improve the quality and efficiency of image generation, allowing for the seamless integration of various image attributes and styles.
@@
These methods offer a one-size-fits-all solution for downstream tasks, which is not suitable for mixed-task scenarios that require personalized services. Moreover, adapter merging methods may encounter issues with parameter interference (Tang et al., 2024), which can severely degrade the performance of the merged model.
2.3. Personalized LoRA serving

### v2

@@ Abstract
Low-Rank Adaptation (LoRA) offers an effective solution for fine-tuning large language models (LLMs). Its modular nature allows the integration of diverse domain-specific LoRAs to enhance LLM capabilities. The emergence of open-source platforms like Huggingface has given rise to a new computational paradigm termed Uploadable Machine Learning (UML). In this framework, edge-side contributors train specialized LoRA adapters and then upload them to a central platform that leverages their plug-and-play nature to provide personalized services for downstream heterogeneous requests. Previous research on LoRA composition has centered on isolated tasks or static adapter selection during training, which does not adequately address the dynamic and expanding pool of LoRAs in UML nor sufficiently meet the needs of downstream heterogeneous requests. To bridge this gap, we propose Retrieval-Augmented Mixture of LoRA Experts (RAMoLE), a retrieve-then-compose framework that adaptively retrieves and composes multiple LoRAs according to the input prompts. RAMoLE contains three main components: firstly, identifying and retrieving LoRAs relevant to the given input through LoraRetriever; secondly, coordinating various dynamically retrieved LoRAs through a carefully designed on-the-fly MoLE mechanism; and thirdly, developing efficient batch inference to accommodate heterogeneous requests. The experimental results indicate that RAMoLE consistently outperforms the baselines, emphasizing its effectiveness and flexible scalability.
Index Terms:
Data mining, text mining, natural language processing, and retrieval models.
I Introduction
@@
Figure 1: Illustration of UML. a) LoRAs from various domains and tasks aimed at enhancing specific capabilities of the LLM can be uploaded to or updated to the LoRA pool. b) The multi-LoRA serving framework aims to leverage the plug-and-play nature of LoRAs to offer comprehensive services. c) The downstream tasks, presented in a mixed-task form, require personalized expert routing.
@@
Recently, leveraging the principles of scaling laws [1], Large Language Models (LLMs) like openAI’s GPT-3.5 [2], GPT-4 [3], Meta’s LLama [4] and many other Foundation Models [5, 6, 7, 8] have achieved notable success across various natural language processing (NLP) tasks [9, 10]. However, these models, primarily trained on diverse internet datasets, sometimes struggle in specialized areas. To address this, numerous studies [11, 12, 9] have explored fine-tuning LLMs using domain-specific data, thereby adapting them to fulfill specific needs. Due to the prohibitively high computation costs for fine-tuning LLMs on specific domains, there is a growing shift towards Parameter-Efficient Fine-Tuning (PEFT) [13, 14, 15], which only updates a small fraction of the model’s parameters or integrates new trainable parameters that augment the model’s capabilities.
@@
As interest grows in open-source communities like Hugging Face [^1] https://huggingface.co/, ModelScope [^2] https://modelscope.cn/home, and Civit AI [^3] https://civitai.com/, there is an increasing trend towards sharing numerous PEFT modules trained on a broad spectrum of domains and tasks across these platforms. Within these PEFT methods, LoRA has become the dominant fine-tuning method [16]. LoRA achieves efficiency by decomposing the weight matrices of a pre-trained model into low-rank factors, allowing it to adapt the model with minimal parameter changes. This approach not only significantly reduces computational load but also maintains or improves model performance through targeted updates. Its modularity and plug-and-play capabilities make it highly adaptable for various tasks and applications. This trend is giving rise to a new decentralized computing paradigm, which we refer to as Uploadable Machine Learning (UML) [17, 16, 18]. As shown in Fig.1, this new computing paradigm involves three components: a) The edge-side domain-specific LoRA contributors. These contributors train plug-and-play, domain-specific LoRA modules in a decentralized manner using local, domain-specific data. The trained LoRA parameters are then uploaded to a centralized service platform. b) Centralized Multi-LoRA Serving platform. The server side runs an LLM equipped with a dynamic pool of LoRAs. The purpose of the LoRA pool is to enhance the LLM’s capabilities across various domains, allowing it to deliver strong performance in a wider range of fields. The LLM leverages the plug-and-play capability of LoRA to dynamically load various LoRAs, allowing it to deliver personalized services tailored to diverse downstream requests. c) Mixed-task Downstream Requests. The downstream requests cover a diverse range of tasks and are in mixed-task form. The heterogeneous nature of these requests requires that the server side utilize the appropriate LoRA for different prompts to deliver personalized services.
@@
Recently, some work has focused on how to compose different LoRAs to enhance the model’s capabilities. Several research has explored the integration of the mixture of expert (MoE; [19, 20]) with LoRAs [21, 11, 22, 23, 24]. However, these methods fix the selection of LoRAs during training, lacking the flexibility to dynamically update and scale in scenarios where the LoRA pool may consistently expand in the UML scenario. This rigidity violates the inherently dynamic and uploadable nature of decentralized training involving domain-specific LoRA experts. LoRAHub [25] and AdapterSoup [26] explore composing LoRAs for specific downstream tasks. However, these two methods provide a one-size-fits-all approach to downstream tasks, failing to accommodate the heterogeneous nature of diverse requests in a UML setting.
In this paper, we introduce a Retrieval-Augmented Mixture of Experts (RAMoLE), a retrieve-then-compose framework designed to exploit the plug-and-play nature of LoRA for UML. Our framework consists of three key components: (1) Input-aware LoRA Retrieval: The first step of our framework is aligning user inputs with the corresponding LoRAs through sentence embeddings and is further refined by an instruction fine-tuning  [27, 28] for effective LoRA retrieval. Through the retriever, we achieve a more flexible LoRA routing mechanism, whose training stage is disentangled from the training and inference of the LLM. (2) On-the-fly Mixture of LoRA Experts: After retrieving the corresponding LoRAs for each sample, we need to finely differentiate and compose these LoRAs for targeted generation. To this end, we propose a novel MoLE mechanism that involves training an additional RouterLoRA, which can be seamlessly integrated into the LLM in an off-the-shelf manner. This mechanism uses an attention mechanism to determine the weights corresponding to the retrieved LoRAs, thereby achieving a more flexible and dynamic MoLE architecture. It is worth noting that vanilla LoRA composition methods, which involve either averaging parameters or outputs, assign uniform weight to all loaded LoRAs. Such approaches may not allocate the most effective weights to the most suitable LoRA, potentially leading to suboptimal outcomes. In contrast, our proposed MoLE mechanism, utilizing a RouterLoRA, finely tunes the allocation of weights to each LoRA, dynamically differentiating each one to enhance overall performance (3) Batch Inference of Multiple LoRAs: Most previous work on the input-adaptive inference of LLMs does not support batch inference [29, 26]. To tackle the challenge of heterogeneous batched requests, we construct a unique LoRA mapping matrix for batch samples. This allows for tailored inferences through efficient matrix multiplication, ensuring each request activates its corresponding LoRAs while maintaining batch processing efficiency.
@@
We note that a shorter conference version of this paper appeared in [17]. Our conference version of the paper did not explicitly define the entire scenario. This manuscript more thoroughly and comprehensively defines the setting as Uploadable Machine Learning. Additionally, the conference version simply applied the same weights for composing the retrieved LoRAs, an approach that may not allocate the most effective weights to the most appropriate LoRA, potentially resulting in suboptimal outcomes. In this manuscript, we introduce a novel on-the-fly MoLE method that employs a RouterLoRA to distinguish between each retrieved LoRA, enhancing performance in the UML scenario. Additionally, to accommodate the new MoLE mechanism, the batch inference mechanism has been adapted to facilitate a more efficient inference process.
Our contribution can be summarized as follows.
@@
II Related Work
II-A Mixture of Experts
@@
The Mixture of Experts (MoE) method combines various specialized sub-modules, guided by a gating network to tailor responses to different input types [19, 20, 30, 31, 32]. Some work [21, 22, 33, 11, 32, 34, 35] focuses on using the MoE method for PEFT to achieve more effective and efficient model fine-tuning. Other work [24, 23] focuses on the use of MoE to post hoc coordinate existing LoRA experts without specifically training the experts’ parameters. Specifically, AdaMix [21] introduces a mixture-of-adaptations approach for fine-tuning models efficiently by combining multiple adaptation methods to reduce parameter overhead. [34] demonstrates the efficacy of combining MoE with lightweight parameter-efficient tuning methods to enhance instruction tuning for large language models, achieving performance comparable to full fine-tuning while significantly reducing computational resource requirements. [22] enhances MoE by developing a highly parameter-efficient model that achieves performance on par with full fine-tuning while significantly reducing parameter usage. SiRA [33] introduces a sparse mixture approach to low-rank adaptation, improving the efficiency and performance of large language models by optimizing sparse computation and reducing overfitting. MoELoRA [11] leverages MoE and LoRA to efficiently fine-tune large language models for diverse medical tasks, addressing the issues of task variety and high computational cost. LoRAMoE [32] integrates MoE and LoRA to enhance large language models’ capability in maintaining and applying world knowledge efficiently, significantly improving model performance while reducing computational requirements. Mixture-of-LoRAs [35] presents a method that leverages a mixture of LoRA modules to enhance the multitask performance of large language models, achieving significant efficiency in fine-tuning across various tasks. These works focus on integrating MoE and LoRA for more efficient fine-tuning of LLMs and reducing computational resources during inference. Our work is orthogonal to these approaches, as we investigate post-hoc LoRA combination, aligning with recent trends in adapting a large LoRA pool for various downstream tasks.
@@
MoLE [24] proposes an inference-time MoE method to enhance the parameter efficiency and flexibility of large language models, allowing for improved fine-tuning and adaptation across diverse tasks without significant computational overhead. SMEAR [23] improves the performance of mixture-of-experts (MoE) models by using a soft merging approach for expert parameters combined with adaptive routing, avoiding the inefficiencies of discrete routing decisions and resulting in better expert specialization and overall model efficiency. PATGOOSE [36] explores a framework for zero-shot learning where a model learns to dynamically route inputs among specialized experts, thereby achieving generalization to unseen tasks by leveraging the expertise of various pre-trained modules, significantly enhancing the model’s adaptability and performance in zero-shot scenarios. [37] presents a modular approach for LLMs by constructing and reusing a library of LoRA modules, enabling efficient fine-tuning and adaptation for multiple tasks, thereby facilitating the creation of versatile and resource-efficient LLMs that can be easily customized for different applications These methods necessitate that the router has encountered all LoRAs during training and requires the selection of LoRAs to be fixed at that time. This limitation prevents these methods from managing the continuously expanding LoRA pool in UML and from routing for unseen LoRAs.
II-B Adapter Merging
@@
In addition to model ensembling through the MoE, there is an increasing focus on aggregating adapters from different domains through the method of Adapter Merging. [38] proposes to compose these parameter-efficient modules through linear arithmetic operations in the weight space to integrate different module capabilities. AdapterSoup [26] aggregates different adapters in the parameter space, allowing large language models to adapt to new domains without additional training. LoRAHub [25] employs random sampling of LoRA parameters from various domains and tasks, followed by black-box optimization to learn the weights of different LoRA parameters without involving model gradient backpropagation. LoRA-Flow [39] dynamically fuses multiple Low-Rank Adaptation (LoRA) modules to enhance the generative capabilities of LLMs, enabling more efficient and effective task-specific fine-tuning without substantial computational overhead. $\pi$-Tuning  [40] considers not only language model experts but also visual model experts. It obtains weights for different LoRAs through task embeddings for downstream tasks, similarly utilizing weighted averaging in the parameter space for targeted model parameter adaptation. LoRA-Composer [41] proposes a framework that utilizes Low-Rank Adaptation to enable multi-concept customization in training-free diffusion models, addressing the challenge of integrating multiple concept LoRAs for complex and varied image generation tasks. [42] explores Multi-LoRA Composition for combining multiple LoRA modules to improve the quality and efficiency of image generation, allowing for the seamless integration of various image attributes and styles.
@@
These methods offer a one-size-fits-all solution for downstream tasks, which is not suitable for mixed-task scenarios that require personalized services. Moreover, adapter merging methods may encounter issues with parameter interference [43], which can severely degrade the performance of the merged model.
II-C Personalized LoRA serving